# Approximation by perfect complexes detects Rouquier dimension

### v1

@@ 1. Introduction
Our work establishes a technical result concerning retracts of good approximations in triangulated categories with a bounded $t$-structure, cf. Lemma 3.2. This result provides a valuable perspective for proving sharp bounds on the Rouquier dimension of triangulated categories of interest in algebraic geometry.
@@
The concept of generation in a triangulated category $\mathcal{T}$ was initially introduced in [BvdB03]. For an object $G$ of $\mathcal{T}$, we denote $\langle G\rangle_{n}$ as the smallest full subcategory of $\mathcal{T}$ generated from $G$ using finite coproducts, shifts, retracts, and at most $n$ cones. The Rouquier dimension of $\mathcal{T}$ is the minimal integer $n$ such that there exists an object $G$ of $\mathcal{T}$ satisfying $\langle G\rangle_{n}=\mathcal{T}$; any object such that $\langle G\rangle_{n}=\mathcal{T}$ for some $n$ is called a strong generator. For additional background, refer to Definition 2.5.
@@
Consider a Noetherian scheme $X$, and let $D^{b}_{\operatorname{coh}}(X)$ denote the derived category of bounded complexes with coherent cohomology. There has been progress made regarding when $D^{b}_{\operatorname{coh}}(X)$ admits a strong generator [Nee21, Aok21, Jat21], as well as on how to identify such objects explicitly [${}^{+}$23, Ola23, Rou08, Orl09, Lan23]. For a smooth variety $X$ over a field, the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is bounded above by $2\dim X$ [Rou08, Ola22]. Initially formulated for smooth quasi-projective varieties in [Orl09], it is conjectured that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ for a smooth variety $X$ coincides with $\dim X$. This expectation has been verified in various instances [Rou08, BF12, Orl09, BFK14, HHL23, Pir23].
@@
Calculating the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ can be challenging. However, when certain mild assumptions on $X$ hold, powerful tools become available. For example, if $X$ is a projective scheme over a commutative Noetherian ring, the converse coghost lemma [OŠ12] offers a valuable method. This lemma facilitates the translation of questions about generation into concerns regarding the vanishing of compositions of specific maps in the bounded derived category. In [OŠ12], the converse coghost lemma was used to show that for any projective scheme $X$ over a commutative Noetherian ring, any strongly generated strictly full triangulated subcategory of $D^{b}_{\operatorname{coh}}(X)$ that contains every perfect complex must coincide with $D^{b}_{\operatorname{coh}}(X)$, see [OŠ12, Theorem 2]. We extend this here to arbitrary Noetherian schemes and additionally prove that if every perfect complex can be built in at most $d$ cones from a given object, then so can $D^{b}_{\operatorname{coh}}(X)$. As discussed in Remark 3.6, every object $E$ in $D^{b}_{\operatorname{coh}}(X)$ has a good approximation by objects in the subcategory $\operatorname{perf}X$, which consists of perfect complexes on $X$. The subsequent result is our first application of Lemma 3.2.
@@
Theorem 1.1 serves as a special case of a broader statement that we establish later regarding triangulated categories with a bounded $t$-structure; refer to Proposition 3.5 for details. For this reason, our result holds more generally for Noetherian algebraic spaces or left/right Noetherian non-commutative rings since approximation by perfect complexes also holds in these contexts. Our proof strategy crucially doesn’t impose any $\operatorname{hom}$-finiteness assumptions on the triangulated category as was required in [OŠ12, Theorem 2], which allows the extension from projective schemes to more general ones can be applied to produce good bounds on Rouquier dimension of singular varieties. For instance, we use it to reprove [BDG17, Corollary 7.2] which says that the bounded derived category of a curve with only nodes and cusps as singularities has Rouquier dimension $\leq 2$, see Corollary 3.11. In what follows is a different flavor of applications of Lemma 3.2.
@@
    If $f\colon U\to X$ is a separated étale morphism of finite type, then every object of $D^{b}_{\operatorname{coh}}(U)$ is a direct summand of $\mathbb{L}f^{\ast}$ of an object of $D^{b}_{\operatorname{coh}}(X)$. In particular, one has
@@
Part (1) was known with $D^{b}_{\operatorname{coh}}$ replaced by $\operatorname{perf}$, see [Bal16, Theorem 3.5]. In the special case that $\mathcal{O}_{X}\to\mathbb{R}f_{\ast}\mathcal{O}_{Y}$ is an isomorphism, part (2) follows from an argument of Kawamata [Kaw22, Lemma 7.4]. A consequence of Theorem 1.2 is demonstrated in Example 3.17, where the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is shown to be no more than twice the Krull dimension of $X$ for (birational) derived splinter varieties over a field. Furthermore, under the assumption that [Orl09, Conjecture 10] holds, the Rouquier dimension of the bounded derived category would become the Krull dimension for such varieties. This observation was previously known only for varieties with rational singularities.
@@
We prove this using a blend of algebraic and geometric arguments, namely Letz’s converse ghost Lemma and Deligne’s formula, cf. Remark 2.10 for further background. As applications, we prove that the Rouquier dimension of the bounded derived category of a Noetherian local ring with a dualizing complex is invariant under (strict) Henselization (Corollary 3.27), and that the bounded derived category of an affine nodal curve is 1 (Corollary 3.28).
@@ Acknowledgements.
The authors would like to thank Amnon Neeman, Srikanth B. Iyengar, Josh Pollitz, Souvik Dey, and Jan Šťovíček for taking a look at an earlier draft of our work. The second author is grateful to the Hausdorff Research Institute for Mathematics for its support and hospitality while this work was being carried out. The recently posted paper [${}^{+}$23] on approximability and bounded $t$-structures touches on similar subject matter but proves a different flavor of results. It would be interesting to find a relation between their work and ours.
@@
  • We will occasionally abuse notation and write $D^{b}_{\operatorname{coh}}(R)$ instead of $D^{b}_{\operatorname{coh}}(\operatorname{Spec}(R))$ for a Noetherian ring $R$. This coincides with the bounded derived category of finitely generated $R$-modules.
@@ 2. Generation
The specific presentation for $(3)$ occurs in [Kra23].
@@
    [Lan23, Example 4.3] Let $\pi\colon\widetilde{X}\to X$ be a proper birational morphism of projective varieties over a field where $\widetilde{X}$ is smooth. If $\mathcal{L}$ is a very ample line bundle on $\widetilde{X}$, then $\mathbb{R}\pi_{\ast}(\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
    [${}^{+}$23, Corollary 3.9] Let $X$ be a quasi-projective variety over a perfect field of nonzero characteristic. Given a very ample line bundle $\mathcal{L}$ on $X$, $F_{\ast}^{e}(\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$ where $e\gg 0$ and $F\colon X\to X$ is the Frobenius morphism.
@@
[…]
@@
This is readily verified from the proofs of Lemma 2.4-2.6 in [Nee21]. Note that strong $\oplus$-generators with bounded coherent cohomology exist in $D_{\operatorname{Qcoh}}(X)$ when $X$ is a quasi-compact separated quasi-excellent scheme of finite Krull dimension, cf. [Aok21, Main Theorem].
@@
[…]
@@
We will call a morphism $f\colon E\to F$ in $\mathcal{T}$ an isomorphism in degree $n$ if $H^{n}(f)\colon H^{n}(E)\to H^{n}(F)$ is an isomorphism in $\mathcal{T}^{\heartsuit}$. The amplitude of a bounded object $E$ of $\mathcal{T}$ is $\infty$- ∞ if the object is zero and
@@ 3. Approximations
Consider a triangulated category $\mathcal{T}$ equipped with a non-degenerate $t$-structure. Let $a,N$ be integers with $N>0$. Choose $E\in\mathcal{T}^{\geq a}$, and assume the following exists:
  • A map $\alpha\colon P\to E$ in $\mathcal{T}$ which is an isomorphism in degrees $\geq a-N$, and
  • A map $F\to F^{\prime}$ in $\mathcal{T}$ which is an isomorphism in degrees $\geq a$ and such that $F^{\prime}\in\mathcal{T}^{\geq a-N}$.
@@
It suffices to show there exist unique dashed arrows which make the diagram above commute. In both cases, this follows directly from Lemma 3.1, where the same argument ensures there exists a unique map $E\to E$ that makes the outer square of the diagram commute. Consequently, we conclude that the composition $E\to F^{\prime}\to E$ is the identity map of $E$. ∎
@@
where $A\to A^{\prime}$ is the projection. This follows from Lemma 3.1. Hence, from this square we obtain the remainder of the desired diagram via the axioms of a triangulated category. It is easy to check that the map $F\to F^{\prime}$ satisfies our claim. This completes the proof. ∎
@@
This follows directly from Proposition 3.5 where we take $\mathcal{P}$ to be $\operatorname{perf}X$. See Remark 3.6 above. ∎
@@
[…]
@@
Following Corollary 3.11, if either $X$ is affine or $X$ is projective over a field $k$ and $Y$ is smooth (for instance if $k$ is perfect), then it is known that $\dim D^{b}_{\operatorname{coh}}(Y)=1$, and so we get $\dim D^{b}_{\operatorname{coh}}(X)\leq 2$. This should be compared to [BD11, Theorem 10], [BDG17, Corollary 7.2], and [BD17, Corollary 3.11].
@@
If $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$, then there exists an $N\geq 0$ such that $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{N}$. Given any object $E$ of $D^{-}_{\operatorname{coh}}(X)$, [Sta23, 0DJN] tells us there exists a distinguished triangle
@@
Let $E$ be an object of $D^{-}_{\operatorname{coh}}(X)$. By the projection formula, we have $\mathbb{R}\pi_{\ast}\mathcal{O}_{Y}\overset{\mathbb{L}}{\otimes}E\cong\mathbb{R}\pi_{\ast}\mathbb{L}\pi^{\ast}E$. Note that $\mathbb{L}\pi^{\ast}E$ is in $D^{-}_{\operatorname{coh}}(Y)$. If $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(Y)$, then Lemma 3.13 tells us there exists an $n\geq 0$ such that $D^{-}_{\operatorname{coh}}(Y)$ is contained in $\overline{\langle G\rangle}_{n}$. Hence, we see that $\mathbb{R}\pi_{\ast}\mathbb{L}\pi^{\ast}E$ is contained in $\overline{\langle\mathbb{R}\pi_{\ast}G\rangle}_{n}$, and so $E$ belongs to $\overline{\langle\mathbb{R}\pi_{\ast}G\rangle}_{nk}$. Remark 2.8 tells us that $E$ belongs to $\langle\mathbb{R}\pi_{\ast}G\rangle_{nk}$. In other words, $D^{b}_{\operatorname{coh}}(X)$ is contained in $\langle\mathbb{R}\pi_{\ast}G\rangle_{nk}$, which completes the proof. ∎
@@
By Lemma 3.15, if $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbb{R}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. Choose a perfect complex $P$ on $X$. If $n$ is the level of $\mathcal{O}_{X}$ with respect to $\mathbb{R}\pi_{\ast}G$ in $D^{b}_{\operatorname{coh}}(X)$, then tensoring with $P$ tells us $P$ belongs to $\langle\mathbb{R}\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P\rangle_{N}$. However, projection formula ensures that $\mathbb{R}\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P$ is isomorphic to $\mathbb{R}\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)$. Suppose the generation time of $G$ in $D^{b}_{\operatorname{coh}}(Y)$ is $g$. Then $G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P$ belongs to $\langle G\rangle_{g+1}$, and so, $\mathbb{R}\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P$ belongs to $\langle\mathbb{R}\pi_{\ast}G\rangle_{g+1}$. Tying this together, we see that $P$ belongs to $\langle\mathbb{R}\pi_{\ast}G\rangle_{(g+1)n}$, and so, the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is bounded above by $(g+1)n-1$. This completes the proof. ∎
@@
Recall that a Noetherian scheme $X$ is a derived splinter (or birational derived splinter) if, for every proper (or birational) surjective morphism $\pi\colon Y\to X$, the natural map $\mathcal{O}_{X}\to\mathbb{R}\pi_{\ast}\mathcal{O}_{Y}$ splits in the derived category $D_{\operatorname{Qcoh}}(X)$. For $X$ a variety over the complex numbers, this is equivalent to $X$ having rational singularities [Kov00, Bha12]. However, for $X$ of prime characteristic, rational singularities are distinct from being a derived splinter, as seen in [Bha12, Example 2.11]. For additional background information, please refer to [Bha12, Lyu22]. If $X$ is (birational) derived splinter variety over a field and $X$ admits a resolution of singularities, Proposition 3.16 asserts that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is at most $2\dim X$. Notably, if [Orl09, Conjecture 10] holds and $X$ admits a resolution of singularities, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is precisely $\dim X$.
@@
By absolute Noetherian approximation [TT, Appendic C] and flat base change, this reduces to the case when $X$ is of finite type over $\mathbb{Z}$. In this case the result follows from [Sta23, Tag 0CRM] via the equivalence [Sta23, Tag 09M5]. ∎
@@
Let $f\colon U\to X$ be a quasi-compact and separated étale morphism of schemes. Then every object $K\in D_{\operatorname{Qcoh}}(X)$ is a direct summand of the object $\mathbb{L}f^{\ast}\mathbb{R}f_{\ast}K$.
@@
(1) Choose an object $E$ in $D^{b}_{\operatorname{coh}}(U)$. By Remark 3.6, there exist a perfect complex $P$ on $U$ and an integer $a$ such that $E=\tau_{\geq a}(P)$. Moreover, Theorem 3.18 guarantees there exists another perfect complex $Q$ on $X$ such that $P$ is a direct summand of $\mathbb{L}f^{\ast}Q$. Since $\tau_{\geq a}$ is a functor, we see that $E=\tau_{\geq a}(P)$ is a direct summand of $\tau_{\geq a}(\mathbb{L}f^{\ast}Q)=\mathbb{L}f^{\ast}(\tau_{\geq a}Q)$, and so we are done.
@@
(3.0.1) $\cdots\to\mathbb{R}g_{\ast}\overline{G}\to\mathbb{R}g_{\ast}\overline{G}\to\cdots\to\mathbb{R}g_{\ast}\overline{G}$
Deligne proves in [Har66, Appendix by P. Deligne] that the assignment $G\mapsto$ the pro-system (3.0.1) is adjoint to the functor $\mathbb{L}f^{\ast}$ in the following precise sense.

### v2

@@ 1. Introduction
Our work establishes a technical result concerning retracts of good approximations in triangulated categories with a bounded $t$-structure, see Lemma 3.2. This result provides a valuable perspective for proving sharp bounds on the Rouquier dimension of triangulated categories of interest in algebraic geometry.
@@
The concept of generation in a triangulated category $\mathcal{T}$ was initially introduced in [BvdB03]. For an object $G$ of $\mathcal{T}$, we denote $\langle G\rangle_{n}$ as the smallest full subcategory of $\mathcal{T}$ generated from $G$ using finite coproducts, shifts, retracts, and at most $n$ cones. The Rouquier dimension of $\mathcal{T}$ is the minimal integer $n$ such that there exists an object $G$ of $\mathcal{T}$ satisfying $\langle G\rangle_{n}=\mathcal{T}$; any object such that $\langle G\rangle_{n}=\mathcal{T}$ for some $n$ is called a strong generator. More generally, an object $G$ is called a classical generator if $\bigcup^{\infty}_{n=0}\langle G\rangle_{n}=\mathcal{T}$. For additional background, refer to Definition 2.5.
@@
Consider a Noetherian scheme $X$, and let $D^{b}_{\operatorname{coh}}(X)$ denote the derived category of bounded complexes with coherent cohomology. There has been progress made regarding when $D^{b}_{\operatorname{coh}}(X)$ admits a strong or classical generators [Nee21, Aok21, Jat21, DLM24a, DLM24b], as well as on how to identify such objects explicitly [BIL+23, Ola23, Rou08, Orl09, Lan24]. The applications of generation for $D^{b}_{\operatorname{coh}}(X)$ have geometrically flavored consequences too. This includes detecting openness of the regular locus for $X$ [DL24a, IT19], as well as characterizations for singularities arising in birational geometry [LV24, LMV25].
For a smooth variety $X$ over a field, the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is bounded above by $2\dim X$ [Rou08, Ola22]. Initially formulated for smooth quasi-projective varieties in [Orl09], it is conjectured that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ for a smooth variety $X$ coincides with $\dim X$. This expectation has been verified in various instances [Rou08, BF12, Orl09, BFK14, HHL23, Pir23].
@@
Calculating the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ can be challenging. However, when certain mild assumptions on $X$ hold, powerful tools become available. For example, if $X$ is a projective scheme over a commutative Noetherian ring, the converse coghost lemma [OŠ12] offers a valuable method. This lemma facilitates the translation of questions about generation into concerns regarding the vanishing of compositions of specific maps in the bounded derived category.
In [OŠ12], the converse coghost lemma was used to show that for any projective scheme $X$ over a commutative Noetherian ring, any strongly generated strictly full triangulated subcategory of $D^{b}_{\operatorname{coh}}(X)$ that contains every perfect complex must coincide with $D^{b}_{\operatorname{coh}}(X)$, see [OŠ12, Theorem 2]. We extend this here to arbitrary Noetherian schemes and additionally prove that if every perfect complex can be built in at most $d$ cones from a given object, then so can $D^{b}_{\operatorname{coh}}(X)$. As shown in [LN07], every object $E$ in $D^{b}_{\operatorname{coh}}(X)$ has a good approximation by objects in the subcategory $\operatorname{perf}X$, which consists of perfect complexes on $X$. The subsequent result is our first application of Lemma 3.2.
@@
Theorem 1.1 serves as a special case of a broader statement that we establish later regarding triangulated categories with a bounded $t$-structure; refer to Proposition 3.5 for details. For this reason, our result holds more generally for Noetherian algebraic spaces or left/right Noetherian non-commutative rings since approximation by perfect complexes also holds in these contexts. Our proof strategy crucially doesn’t impose any $\operatorname{hom}$-finiteness assumptions on the triangulated category as was required in [OŠ12, Theorem 2], which allows the extension from projective schemes to more general ones. Theorem 1.1 can be applied to produce good bounds on Rouquier dimension of singular varieties. For instance, we use it to reprove [BDG17, Corollary 7.2] which says that the bounded derived category of a curve with only nodes and cusps as singularities has Rouquier dimension $\leq 2$, see Corollary 3.10. In what follows is a different flavor of applications of Lemma 3.2.
@@
    If $f\colon U\to X$ is a quasi-compact separated étale morphism [^1] These hypotheses will ensure that $U$ is Noetherian., then every object of $D^{b}_{\operatorname{coh}}(U)$ is a direct summand of $\mathbb{L}f^{\ast}$ of an object of $D^{b}_{\operatorname{coh}}(X)$. In particular, one has
@@
Part (1) was known with $D^{b}_{\operatorname{coh}}$ replaced by $\operatorname{perf}$, see [Bal16, Theorems 3.5 and 4.2]. In the special case that $\mathcal{O}_{X}\to\mathbb{R}f_{\ast}\mathcal{O}_{Y}$ is an isomorphism, part (2) follows from an argument of Kawamata [Kaw22, Lemma 7.4]. A consequence of Theorem 1.2 is demonstrated in Example 3.16, where the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is shown to be no more than twice the Krull dimension of $X$ for (birational) derived splinter varieties over a field. Furthermore, under the assumption that [Orl09, Conjecture 10] holds, the Rouquier dimension of the bounded derived category would become the Krull dimension for such varieties. This observation was previously known only for varieties with rational singularities.
@@
We prove this using a blend of algebraic and geometric arguments, namely Letz’s converse ghost Lemma and Deligne’s formula, see Remark 2.10 for further background. As applications, we prove that the Rouquier dimension of the bounded derived category of a Noetherian local ring with a dualizing complex is invariant under (strict) Henselization (Corollary 3.26), and that the bounded derived category of an affine nodal curve is 1 (Corollary 3.27).
@@ Acknowledgements.
The authors would like to thank Amnon Neeman, Srikanth B. Iyengar, Josh Pollitz, Souvik Dey, and Jan Šťovíček for taking a look at an earlier draft of our work. The second author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813 and by the European Union (ERC, ZETA-FM, 8641450). He is grateful to the Hausdorff Research Institute for Mathematics for its support and hospitality while this work was being carried out. The recently posted paper [BCM+23] on approximability and bounded $t$-structures touches on similar subject matter but proves a different flavor of results. It would be interesting to find a relation between their work and ours. Both authors thank the anonymous referee for help suggestions on improvements to our work.
@@
  • We will occasionally abuse notation and write $D^{b}_{\operatorname{coh}}(R)$ instead of $D^{b}_{\operatorname{coh}}(\operatorname{Spec}(R))$ for a Noetherian ring $R$, which coincides with the bounded derived category of finitely generated $R$-modules.
@@ 2. Generation
The specific presentation for $(3)$ occurs in [Kra24].
@@
    [Lan24, Example 4.3] Let $\pi\colon\widetilde{X}\to X$ be a proper birational morphism of projective varieties over a field where $\widetilde{X}$ is smooth. If $\mathcal{L}$ is a very ample line bundle on $\widetilde{X}$, then $\mathbb{R}\pi_{\ast}(\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$.
@@
    [BIL+23, Corollary 3.9] Let $X$ be a quasi-projective variety over a perfect field of nonzero characteristic. Given a very ample line bundle $\mathcal{L}$ on $X$, $F_{\ast}^{e}(\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$ where $e\gg 0$ and $F\colon X\to X$ is the Frobenius morphism.
@@
([BvdB03, §2.2§2.2$\S$ 2.2§ 2.2], [Nee21, §$\S 1$], [Rou08, §3.3§3.3$\S$ 3.3§ 3.3]) Let $\mathcal{S}$ be a subcategory of $\mathcal{T}$, and assume that $\mathcal{T}$ admits all small coproducts.
@@
This is readily verified from the proofs of Lemma 2.4-2.6 in [Nee21]. The special case where $G$ is in $\operatorname{perf}X$ is [BvdB03, Proposition 2.2.4]. Note that strong $\oplus$-generators with bounded coherent cohomology exist in $D_{\operatorname{Qcoh}}(X)$ when $X$ is a quasi-compact separated quasi-excellent scheme of finite Krull dimension, see [Aok21, Main Theorem].
@@
These inequalities are called the coghost and ghost lemmas. If $\mathcal{T}=D^{b}_{\operatorname{coh}}(X)$ where $X$ is a projective scheme over a commutative Noetherian ring, then $\operatorname{cogin}^{G}_{\mathcal{T}}(E)=\operatorname{level}^{G}_{\mathcal{T}}(E)$ by [OŠ12, Theorem 4]. This (more precisely, the inequality $\geq$) is sometimes called the converse coghost lemma. In the case $X$ admits a dualizing complex, one can deduce that also $\operatorname{gin}^{G}_{\mathcal{T}}(E)=\operatorname{level}^{G}_{\mathcal{T}}(E)$. See [Let21, §$\S 2.13$] for the special case when $X$ is an affine Noetherian scheme with a dualizing complex, which is all we need in this paper. We refer to this (more precisely, the inequality $\geq$) as the converse ghost lemma.
@@
We will call a morphism $f\colon E\to F$ in $\mathcal{T}$ an isomorphism in degree $n$ if one has $H^{n}(f)\colon H^{n}(E)\to H^{n}(F)$ is an isomorphism in $\mathcal{T}^{\heartsuit}$. The amplitude of a bounded object $E$ of $\mathcal{T}$ is $\infty$- ∞ if the object is zero and
@@ 3. Approximations
Let $\mathcal{T}$ be a triangulated category equipped with a non-degenerate $t$-structure. Let $a,N$ be integers with $N>0$. Suppose given two morphisms $P\to E$ and $F\to F^{\prime}$ in $\mathcal{T}$. Assume:
  • $P\to E$ is an isomorphism in degrees $\geq a-N$ and $E$ is in $\mathcal{T}^{\geq a}$,
  • $F\to F^{\prime}$ is an isomorphism in degrees $\geq a$ and $F^{\prime}\in\mathcal{T}^{\geq a-N}$.
@@
It follows from Lemma 3.1 that there exist unique dashed arrows making the diagram commute: The morphisms
$\operatorname{Hom}(E,F^{\prime})\to\operatorname{Hom}(P,F^{\prime}),\hskip 20.00003pt\operatorname{Hom}(F^{\prime},E)\to\operatorname{Hom}(F,E)$
are both isomorphisms. The same argument ensures there exists a unique map $E\to E$ that makes the outer square of the diagram commute. Consequently, we conclude that the composition $E\to F^{\prime}\to E$ is the identity map of $E$. ∎
@@
where $A\to A^{\prime}$ is the projection. This follows from Lemma 3.1. Hence, from this square we obtain the remainder of the desired diagram via the axioms of a triangulated category. Since $A\to A^{\prime}$ is an isomorphism in degrees $>a$ and $B\to B^{\prime}$ is an isomorphism in degrees $\geq a$, we see that $F\to F^{\prime}$ is an isomorphism in degrees $\geq a$, as needed. ∎
@@
This follows directly from Proposition 3.5 where we take $\mathcal{P}$ to be $\operatorname{perf}X$. See [LN07, Theorem 4.1]. ∎
@@
Recall that the delta invariant of a one-dimensional reduced Nagata Noetherian ring $A$ is the length of the $A$-module $B/A$, where $B$ is the normalisation of $A$. Thus, $A$ has delta invariant one if, and only if, $B/A\cong A/\mathfrak{m}_{A}$ as $A$-modules. The delta invariant is one for the local ring of a curve at a nodal or cuspidal singularity, see [Har83, Ex. §$\S 3.8.1$].
@@
Following Corollary 3.10, if either $X$ is affine or $X$ is projective over a field $k$ and $Y$ is smooth (for instance if $k$ is perfect), then it is known that $\dim D^{b}_{\operatorname{coh}}(Y)=1$ [Orl09], and so we get $\dim D^{b}_{\operatorname{coh}}(X)\leq 2$. This should be compared to [BD11, Theorem 10], [BDG17, Corollary 7.2], and [BD17, Corollary 3.11].
@@
If $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$, then there exists an $N\geq 0$ such that $D^{b}_{\operatorname{coh}}(X)=\langle G\rangle_{N}$. Given any object $E$ of $D^{-}_{\operatorname{coh}}(X)$, [Sta23, 0DJN] tells us $E$ is a homotopy colimit of perfect objects. This gives us a distinguished triangle:
@@
The following lemma has significant generalizations which appear elsewhere, see [DL24b, Theorem 3.11] and for a noncommutative version refer to [DLM24b, Theorem 3.12]. But we record it anyways for its relevance to our work in action.
@@
Let $E$ be an object of $D^{-}_{\operatorname{coh}}(X)$. By the projection formula [Sta23, Tag 08EU], we have $\mathbb{R}\pi_{\ast}\mathcal{O}_{Y}\overset{\mathbb{L}}{\otimes}E\cong\mathbb{R}\pi_{\ast}\mathbb{L}\pi^{\ast}E$. Note that $\mathbb{L}\pi^{\ast}E$ is in $D^{-}_{\operatorname{coh}}(Y)$. If $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(Y)$, then Lemma 3.12 tells us there exists an $n\geq 0$ such that $D^{-}_{\operatorname{coh}}(Y)$ is contained in $\overline{\langle G\rangle}_{n}$. Hence, we see that $\mathbb{R}\pi_{\ast}\mathbb{L}\pi^{\ast}E$ is contained in $\overline{\langle\mathbb{R}\pi_{\ast}G\rangle}_{n}$, and so after taking more cones, we see $E$ belongs to $\overline{\langle\mathbb{R}\pi_{\ast}G\rangle}_{nk}$. Remark 2.8 tells us that $E$ belongs to $\langle\mathbb{R}\pi_{\ast}G\rangle_{nk}$. In other words, $D^{b}_{\operatorname{coh}}(X)$ is contained in $\langle\mathbb{R}\pi_{\ast}G\rangle_{nk}$, which completes the proof. ∎
@@
By Lemma 3.14, if $G$ is a strong generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbb{R}\pi_{\ast}G$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$. Choose a perfect complex $P$ on $X$. If $n$ is the level of $\mathcal{O}_{X}$ with respect to $\mathbb{R}\pi_{\ast}G$ in $D^{b}_{\operatorname{coh}}(X)$, then tensoring with $P$ tells us $P$ belongs to $\langle\mathbb{R}\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P\rangle_{n}$. However, projection formula ensures that $\mathbb{R}\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P$ is isomorphic to $\mathbb{R}\pi_{\ast}(G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P)$. Suppose the generation time of $G$ in $D^{b}_{\operatorname{coh}}(Y)$ is $g$. Then $G\overset{\mathbb{L}}{\otimes}\mathbb{L}\pi^{\ast}P$ belongs to $\langle G\rangle_{g+1}$, and so, $\mathbb{R}\pi_{\ast}G\overset{\mathbb{L}}{\otimes}P$ belongs to $\langle\mathbb{R}\pi_{\ast}G\rangle_{g+1}$. Tying this together, we see that $P$ belongs to $\langle\mathbb{R}\pi_{\ast}G\rangle_{(g+1)n}$, and so, the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is bounded above by $(g+1)n-1$ by Theorem 1.1. This completes the proof. ∎
@@
Recall that a Noetherian scheme $X$ is a derived splinter (respectively birational derived splinter) if, for every proper (respectively birational) surjective morphism $\pi\colon Y\to X$, the natural map $\mathcal{O}_{X}\to\mathbb{R}\pi_{\ast}\mathcal{O}_{Y}$ splits in the derived category $D_{\operatorname{Qcoh}}(X)$. For $X$ a variety over the complex numbers, this is equivalent to $X$ having rational singularities [Kov00, Bha12]. However, for $X$ of prime characteristic, rational singularities are distinct from being a derived splinter, as seen in [Bha12, Example 2.11]. For additional background information, please refer to [Bha12, Lyu22]. If $X$ is (birational) derived splinter variety over a field and $X$ admits a resolution of singularities $\widetilde{X}\to X$, Proposition 3.15 asserts that the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is at most the Rouquier dimension of $D^{b}_{\operatorname{coh}}(\widetilde{X})$, which is at most $2\dim X$ by [Rou08, Proposition 7.9]. Notably, if [Orl09, Conjecture 10] holds and $X$ admits a resolution of singularities, then the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ is precisely $\dim X$. On a different topic, see [LV24] for a triangulated characterization of derived splinters in terms of generation.
@@
By absolute Noetherian approximation [TT, Appendix C] and flat base change, this reduces to the case when $X$ is of finite type over $\mathbb{Z}$. In this case the result follows from [Sta23, Tag 0CRM] via the equivalence [Sta23, Tag 09M5]. ∎
@@
([Bal16, proof of Theorem 3.5]) Let $f\colon U\to X$ be a quasi-compact and separated étale morphism of schemes. Then every object $K\in D_{\operatorname{Qcoh}}(X)$ is a direct summand of the object $\mathbb{L}f^{\ast}\mathbb{R}f_{\ast}K$.
@@
(1) Choose an object $E$ in $D^{b}_{\operatorname{coh}}(U)$. By [LN07, Theorem 4.1], there exist a perfect complex $P$ on $U$ and an integer $a$ such that $E=\tau_{\geq a}(P)$. Moreover, Theorem 3.17 guarantees there exists another perfect complex $Q$ on $X$ such that $P$ is a direct summand of $\mathbb{L}f^{\ast}Q$. Since $\tau_{\geq a}$ is a functor, we see that $E=\tau_{\geq a}(P)$ is a direct summand of $\tau_{\geq a}(\mathbb{L}f^{\ast}Q)=\mathbb{L}f^{\ast}(\tau_{\geq a}Q)$, and so we are done.
@@
(3.1) $\cdots\to\mathbb{R}g_{\ast}\overline{G}\to\mathbb{R}g_{\ast}\overline{G}\to\cdots\to\mathbb{R}g_{\ast}\overline{G}$
Deligne proves in [Har66, Appendix by P. Deligne] that the assignment $G\mapsto$ the pro-system (3.1) is adjoint to the functor $\mathbb{L}f^{\ast}$ in the following precise sense.